OpenStudy (yrivers36):

The grams of carbohydrate in one apple if the apple has no fat and no protein and provides 72 kcal of energy. What ate the kilocalories or kJ?

OpenStudy (aaronq):

the question is incomplete

OpenStudy (yrivers36):

Carbs 4 kcal/g 17 kJ/g Fat 9 kcal/g 38 kJ/g Protein 4 kcal/g 17 kJ/g Using the energy values for foods (above) determine each of the following: The grams of carbohydrate in one apple if the apple has no fat and no protein and provides 72 kcal of energy.

OpenStudy (aaronq):

so you can use dimensional analysis, you need a value in grams, and you have "kcal/g" and "kcal". You need to cancel out "kcal" and have "g" at the top (numerator). \(\dfrac{kcal}{kcal/g}=\dfrac{\cancel{kcal}}{\cancel{kcal}}*g=g\) so divide the energy by the values given in the table. Mass of apple =\(\dfrac{ 72 kcal}{ 4 kcal/g}=18~g\)

OpenStudy (yrivers36):

Thank you so much!